The following information was collected for the firm you are trying to value: Item 2020 ($ M) 2019 ($M) Property, plant, and equipment 15,905 16,126 Intangibles 25,764 27,222 Goodwill 21,629 21,832 Current assets 60,210 55,744 Current liabilities 27,747 25,031 Net earnings before taxes 19,196 20,563 Taxes 4,806 7,548 Depreciation 3,746 3,895 Net debt repaid 1,379 850 Compute the FCFTE for the firm for the year 2020.

Use the information for the FCFTE in 2020. In addition,  If the short-term growth rate (for next five years) is estimated to be 6% and the long-term growth rate beyond that is estimated to be 3%, compute the FCF for your firm for the years 2021 – 2026. Enter your answer as a string of numbers. (For example, if you think the answer is $12,345.67 then enter 12345.67
Answer #1
2020 2019
PP&E 15905 16126
Capex = Change in Net PP&E + Depreciation
Capex = (15905-16126)+3746
Capex = 3525
Intangible 25764 27222
Goodwill 21629 21832
Intangible & Goodwill 47393 49054
Amortization = 49054 - 47393
Amortization = 1661
Current Assets 60210 55744
Current Liabilities 27747 25031
Working capital (Current assets - Current liabilities) 32463 30713
Change in working capital = 32463 - 30713
Change in working capital = 1750
Net earning before taxes 19196 20563
Taxes 4806 7548
Net earning after taxes 14390 13015
Depreciation 3746 3895
Net debt repaid 1379 850
FCFTE = Net income + Depreciation & Amortisation - Change in working capital - Capex + Net borrowing
FCFTE = 3746 + 14390 + 1661 - 1750 - 3525 + 1379
FCFTE = 15901